A normal population has a mean of 52 and a standard deviation of 14. You select a random sample of 37. Use Appendix B.1 for the z-values.

Compute the probability that the sample mean is: (Round the final answers to 4 decimal places.)

a. Greater than 55.

Probability


b. Less than 48.

Probability

c. Between 48 and 55.

Probability

Answer by Boreal(15235)

greater than 55, z=(x bar-mean)/sigma/sqrt(n)
so z>(3/14/sqrt(37))= 3sqrt(37)/14=1.30
that probability is 0.0968
less than 48 would be z< -4/14/sqrt(37)=-4 sqrt(37)/14=-1.74
That probability is 0.0409
Between 48 and 55 is everything else or 1-0.1377=0.8623